# Topic-FlipRAG: Topic-Orientated Adversarial Opinion Manipulation Attacks to Retrieval-Augmented Generation Models

### v2

@@ 6.2. RQ2: To what extent does Topic-FlipRAG affect the answers generated by the target RAG systems?
6.3. RQ3: Does topic-oriented opinion manipulation significantly impact users’ perceptions of controversial topics?
@@
Figure 4. Empirical comparative experiments of adversarial opinion manipulation attacks on user cognition.
@@
To explore the effect of opinion manipulation on users’ perceptions of controversial topics, we conducted a user study involving 54 college students (29 males, 25 females) to assess how opinion manipulation influences perceptions of controversial topics. Participants were split into two equal groups (Group-clean and Group-poisoned, 27 members each), remaining unaware of the study’s true purpose and operating independently. A simple QA service based on RAG was implemented, and Topic-FlipRAG was employed to poison documents with specific stances on two designated controversial topics. For one topic, supporting documents were poisoned; for the other, opposing documents were poisoned. Each participant interacted with the QA system over three rounds per topic, mirroring real-world usage patterns, and subsequently rated their stance on a 7-point Likert scale (normalized to 0 for strong opposition, 1 for strong support).
@@
Figure 4 illustrates the significant shifts in user opinions resulting from Topic-FlipRAG. For Topic 1 (Should People Become Vegetarian?), the manipulated stance was PRO. Group-clean exhibited a dispersed distribution with an average score of 0.45, reflecting diverse views on this controversial issue. By contrast, Group-poisoned recorded a notably higher mean score of 0.61, indicating successful manipulation toward pro-vegetarian sentiment. For Topic 2 (Should Humans Colonize Space?), the stance was CON. Although Group-clean originally showed strong support (average score 0.76), Group-poisoned achieved a reduction to 0.59, effectively nudging opinions toward neutrality.
These findings underscore the potency of Topic-FlipRAG in altering user perceptions via RAG, whether opinions are initially neutral or polarized. The consistent polarity shifts across both topics reveal the potential risks posed by adversarial attacks in shaping public discourse on contentious issues.
@@
Table 4. Ablation study (top3-ori, top3-att, and top3-v are reported in %) on the PROCON dataset. w/o denotes “without”.
@@ 6.4. Ablation Study
Adv-trigger and know-attack complement each other by leveraging distinct strengths to enhance the overall effectiveness of the Topic-FlipRAG framework. Adv-trigger, through its gradient-based generation process, directly optimizes the retriever mechanism, resulting in significant ranking improvements for adversarially targeted documents. In contrast, know-attack, while less impactful on ranking optimization, utilizes general semantic knowledge to achieve precise and consistent polarity manipulation in LLM outputs, reinforcing the adversarial influence on generated responses. This synergy allows for more comprehensive and precise control over the RAG framework’s behavior, thereby maximizing the overall efficacy of the attack strategy.
Impact of Rewarding Function in Topic-FlipRAG. The rewarding function in Topic-FlipRAG dynamically adjusts the augment factor $t$ in the know-attack process (as described in Algorithm 1), controlling the aggressiveness of edits in each iteration. By leveraging semantic similarity and edit distance filtering, the rewarding function ensures an adaptive improvement strategy, enhancing the overall performance.
As shown in Table 5, the rewarding function significantly impacts the effectiveness of know-attack in IR manipulation. Compared to fixed augment factor $t$ values, the dynamic adjustment achieves higher attack breadth (evidenced by increased RASR) and depth (higher top3-v scores). Table 6 further highlights its influence on RAG opinion manipulation, where the dynamic approach yields higher ASV and $\Delta$ ASV compared to fixed $t$ settings. Notably, the rewarding function’s influence on RAG output polarity manipulation surpasses its impact on IR manipulation. This can be attributed to its role in improving the quality of generated $\text{doc}_{\text{know}}$, where quality encompasses both enhanced ranking boosts (related to IR manipulation) and stronger polarity reinforcement. These findings underscore the rewarding function’s critical contribution to optimizing the know-attack strategy for both ranking and opinion manipulation objectives.
Table 5. Comparative analyses of rewarding function and fixed augment factor $t$ on IR manipulation attacks (%).
@@
Table 6. Comparative analyses of rewarding function and fixed augment factor $t$ on RAG opinion manipulation.
@@
$\Delta$ ASV 0.37 0.31 0.29 0.32
Impact of Polarity Control in Topic-FlipRAG. Figure 5 illustrates the effect of the Polarity Control component on the sentiment of adversarial documents ($\text{doc}_{\text{adv}}$) when the target polarity is set to CON. Negative sentiment scores represent negative polarity, while positive scores indicate positive polarity. (For PRO, see Figure 10 in the Appendix.) The results show that, without Polarity Control, $\text{doc}_{\text{adv}}$ gravitates toward a neutral polarity. With Polarity Control enabled, Topic-FlipRAG effectively aligns $\text{doc}_{\text{adv}}$ with the desired target polarity (CON).To further investigate the influence of attacks on document sentiment, we present measurement results in Figure 12 and Figure 13 in the Appendix.
To further investigate how Polarity Control influences RAG opinion manipulation, we examine not only the sentiment but also the stance of $\text{doc}_{\text{adv}}$ (i.e., its degree of alignment with the target polarity). For both PRO and CON settings, we randomly selected five topics and measured stance (normalized from $-1$ to $1$, where $-1$ signifies strong opposition and $1$ signifies strong support). As shown in Figure 11 (in the Appendix), without Polarity Control, the stance of $\text{doc}_{\text{adv}}$ shifts closer to neutrality, weakening its polarity. In contrast, incorporating Polarity Control preserves stance, enabling more effective manipulation of the LLM’s final outputs.
Figure 5. Impact of Polarity Control module on documents sentiment. w/o denotes “without”. (target polarity:CON).
The notable deviation in $\text{doc}_{\text{adv}}$ polarity without Polarity Control arises from the nature of know-attack: while it leverages the internal knowledge and reasoning capabilities of LLMs, it also inherits inherent polarity biases. Absent strict polarity enforcement during each editing step, the generated $\text{doc}_{\text{adv}}$ tends to drift from the intended polarity. By enforcing the desired polarity at every phase of know-attack, the Polarity Control component supports a more consistent and effective manipulation of RAG polarity outcomes.
Figure 6. Log perplexity (PPL) distributions on the PROCON dataset.
Potential Effectiveness by Altering Poisoning Target. Table 7 illustrates the impact of different poisoned methods on IR and RAG manipulation results. In our prior experiments, we selected the least relevant five documents (last5) as the poisoning targets to evaluate Topic-FlipRAG’s performance in a challenging manipulation scenario. This design aimed to test the robustness of the attack under more constrained conditions. However, in real-world attack scenarios, an attacker might reasonably choose the topically most relevant documents (top5) as poisoned targets to achieve more effective opinion manipulation.
The experimental results show that poisoning the top5 documents yields improved manipulation performance, as evidenced by higher top3-v and ASV scores. Specifically, poisoning the top5 documents achieves a top3-v score of 0.37 and an ASV score of 0.71, both surpassing the performance observed with the last5 method. This indicates that Topic-FlipedRAG could cause even greater harm if deployed in real-world scenarios, where attackers have the flexibility to select highly relevant documents as targets. These findings underscore the pressing need to develop robust defenses against such advanced adversarial strategies.
Table 7. Performance comparison of poisoning target on the PROCON dataset.(top3-ori, top3-att, top3-v are reported in %.)
@@
6.5. RQ4: How robust does Topic-FlipRAG against exisiting mitigation mechanism?
Mitigation by perplexity. Perplexity (PPL) is widely used to measure the quality of texts, which is also utilized to defend against attacks to LLMs and IR model(Jain et al., 2023; Gonen et al., 2022; Liu et al., 2022). A large perplexity of a text means it is of low quality. We utilize perplexity to detect malicious texts.
Figure 6 shows the log perplexity (PPL) distributions evaluated using GPT-2 (Radford et al., 2019) on PROCON dataset documents manipulated by various attack methods. The results reveal the limitations of perplexity-based defenses in distinguishing poisoned documents from clean ones, as significant overlaps in PPL distributions persist. This overlap makes it difficult to set effective thresholds: strict thresholds lead to high false positive rates, while loose thresholds fail to detect attacks, undermining the defense’s efficacy.
Specifically,Collision exhibits the largest deviation in PPL distribution, making it relatively easier to detect. In contrast, Topic-FlipRAG and PAT produce poisoned documents with distributions closely resembling the original dataset, posing significant challenges for detection. PoisonedRAG, which lacks gradient-guided optimization, shows almost no shift in PPL, further highlighting the insufficiency of perplexity-based filtering for nuanced manipulations.
Figure 7. Attack performance of different random mask rate.
@@ 6.5. RQ4: How robust does Topic-FlipRAG against exisiting mitigation mechanism?
Mitigation by Random Mask. Random Mask is a robustness enhancement method that randomly masks a proportion of tokens with a placeholder token $m$ during embedding (Zeng et al., 2023). By averaging predictions over multiple masked input versions, it reduces sensitivity to small perturbations and enhances model stability. In our implementation, we averaged outputs over three masked copies of each input to assess performance under varying mask rates.
@@
Figure 7 presents experimental results on the PROCON-Dataset targeting the CON stance. When the mask rate is below 0.3, Topic-FlipRAG achieves notably higher manipulation metrics (RASR ¿ 55%, top3-v ¿ 20%) compared to other baselines, underscoring its robust manipulation capability. As the mask rate increases, the effectiveness of all methods declines; however, Topic-FlipRAG still maintains the highest attack success among the baselines. Nonetheless, excessively large mask rates also degrade normal RAG performance, revealing the trade-off between heightened defense and overall system functionality. Consequently, Random Mask alone is insufficient against sophisticated attacks like Topic-FlipRAG, indicating that additional measures are required to strengthen system robustness without undermining functionality.
@@
Mitigation by reranking. We investigate a reranking mechanism within the RAG framework to defend against adversarial attacks. This approach uses a different IR model to refine and reorder initial retrieval results, thus diminishing attack efficacy. The defense rationale is twofold: (1) existing attacks typically rely on simulating the target IR model (Cho et al., 2024; Liu et al., 2022, 2024), and introducing a distinct reranker weakens this dependency; (2) varying IR models exhibit different robustness levels, and reranking adds resilience.
@@
Figure 8. Comparative analysis of the performance metrics (RASR and top3-v) across different baselines before and after reranking mitigation.
7. Conclusion

### v3

@@ 6.2 RQ2: To what extent does Topic-FlipRAG affect the answers generated by the target RAG systems?
6.3 RQ3: Does topic-oriented opinion manipulation significantly impact users’ perceptions of controversial topics?
@@
Figure 3: Empirical comparative experiments of adversarial opinion manipulation attacks on user cognition.
@@
To explore the effect of opinion manipulation on users’ perceptions of controversial topics, we conducted a user study involving 54 college students to assess how opinion manipulation influences perceptions of controversial topics. Participants were split into two equal groups (Group-clean and Group-poisoned, 27 members each), remaining unaware of the study’s true purpose and operating independently. A simple QA service based on RAG was implemented, and Topic-FlipRAG was employed to poison documents with specific stances on two designated controversial topics. For one topic, supporting documents were poisoned; for the other, opposing documents were poisoned. Each participant interacted with the QA system over three rounds per topic, mirroring real-world usage patterns, and subsequently rated their stance on a 7-point Likert scale (normalized to 0 for strong opposition, 1 for strong support). Full user experimental details are provided in the Appendix A.3.
@@
Figure 3 illustrates the significant shifts in user opinions resulting from Topic-FlipRAG. For Topic 1 (Should People Become Vegetarian?), the manipulated stance was PRO. Group-clean exhibited a dispersed distribution with an average score of 0.45, reflecting diverse views on this controversial issue. By contrast, Group-poisoned recorded a notably higher mean score of 0.61, indicating successful manipulation toward pro-vegetarian sentiment. For Topic 2 (Should Humans Colonize Space?), the stance was CON. Although Group-clean originally showed strong support (average score 0.76), Group-poisoned achieved a reduction to 0.59, effectively nudging opinions toward neutrality. These results underscore Topic-FlipRAG’s potency in shifting both neutral and polarized views, highlighting the potential risks of adversarial opinion manipulation in shaping public discourse.
@@
Table 4: Ablation study (top3-ori, top3-att, and top3-v are reported in %) on the PROCON dataset. w/o denotes “without”.
@@ 6.4 Ablation Study
Adv-trigger and know-attack complement each other by leveraging distinct strengths to enhance the overall effectiveness of the Topic-FlipRAG framework. Adv-trigger, through its gradient-based generation process, directly optimizes the retriever mechanism, resulting in significant ranking improvements for adversarially targeted documents. In contrast, know-attack, while comparatively less impactful on ranking optimization, utilizes general semantic knowledge to achieve precise and consistent polarity manipulation in LLM outputs, reinforcing the adversarial influence on generated responses. This synergy allows for more comprehensive and precise control over the RAG framework’s behavior, thereby maximizing the overall efficacy of the attack strategy.
Impact of Rewarding Function in Topic-FlipRAG. The rewarding function in Topic-FlipRAG dynamically adjusts the augment factor $t$ in the know-attack process, controlling the aggressiveness of edits in each iteration. By leveraging semantic similarity and edit distance filtering, the rewarding function ensures an adaptive improvement strategy, enhancing the overall performance.
[…]
Table 5: Comparative analyses of rewarding function and fixed augment factor $t$ on IR manipulation attacks (%).
@@
Table 6: Comparative analyses of rewarding function and fixed augment factor $t$ on RAG opinion manipulation.
@@
[…]
Impact of Polarity Control in Topic-FlipRAG. Figure 4 illustrates the effect of the Polarity Control component on the sentiment of the $\text{doc}_{adv}$ when the target polarity is set to CON. Negative sentiment scores represent negative polarity, while positive scores indicate positive polarity. The results show that, without Polarity Control, the $\text{doc}_{adv}$ gravitates toward a neutral polarity. With Polarity Control enabled, Topic-FlipRAG effectively aligns $\text{doc}_{adv}$ with the desired target polarity CON).
The notable deviation in the $\text{doc}_{adv}$ polarity without Polarity Control arises from the nature of know-attack: while it leverages the internal knowledge and reasoning capabilities of LLMs, it also inherits inherent polarity biases. Absent strict polarity enforcement during each editing step, the generated $\text{doc}_{adv}$ tends to drift from the intended polarity. By enforcing the desired polarity at every phase of know-attack, the Polarity Control component supports a more consistent and effective manipulation of RAG polarity outcomes.
Figure 4: Impact of Polarity Control module on documents sentiment. w/o denotes “without”. (target polarity:CON).
Table 7: Performance comparison of poisoning target on the PROCON dataset.(top3-ori, top3-att, top3-v are reported in %.)
@@
Figure 5: Impact of the number of retrieved documents $K$ (left part) and the iteration number NN (right part) on different performance metrics of Topic-FlipRAG on the PROCON dataset.
Potential Effectiveness by Altering Poisoning Target. Table 7 illustrates the impact of different poisoned methods on IR and RAG manipulation results.In our prior experiments on the PROCON dataset, we intentionally selected the least relevant five documents (last5) as poisoning targets to assess Topic-FlipRAG’s effectiveness under a challenging manipulation scenario. This design aimed to test the robustness of the attack under more constrained conditions. However, in real-world attack scenarios, an attacker might reasonably choose the topically most relevant documents (top5) as poisoned targets to achieve more effective opinion manipulation.
The experimental results show that poisoning the top5 documents yields improved manipulation performance, as evidenced by higher top3-v and ASV scores. Specifically, poisoning the top5 documents achieves a top3-v score of 0.37 and an ASV score of 0.71, both surpassing the performance observed with the last5 method. These results suggest that Topic-FlipRAG could pose even greater risks in real-world settings, where attackers can freely select highly relevant targets, highlighting the urgent need for robust defenses against such advanced adversarial strategies.
6.5 Hyper-parameter Analysis
We analyze four key hyper-parameters of our method: the editing distance $\epsilon$, the semantic similarity threshold $\lambda$, the iteration number $N$, and the number of retrieved documents $K$. Due to space limitations, the analyses of $\epsilon$ and $\lambda$ are demonstrated in detail in Appendix A.2. We do not discuss the hyper-parameters used in adversarial trigger generation, as they have already been systematically studied and validated in prior ranking attack research [song2020adversarial, liu2022order].
Number of retrieved documents $K$. This hyper-parameter controls how many candidate documents are retrieved for LLM generation. As $K$ varies, we replace the fixed top3-v metric with a dynamic topK-v to reflect stance manipulation across the top-$K$ candidates. RASR is also redefined accordingly: an attack is considered successful if the proportion of target-label documents increases within the top-$K$ results.
Our results (Figure 5) show that RASR increases with $K$, suggesting that when the retriever returns more documents, the chance of adversarial content appearing in the top-$K$ candidates rises, thus increasing the system’s vulnerability to manipulation. However, both topK-v and ASV exhibit a downward trend as $K$ increases. This indicates a dilution effect: while manipulated documents are more likely to be included, their influence on the final answer is weakened due to the expanded pool of candidates. To achieve a balanced evaluation that jointly accounts for both RASR and ASV, we set $K=3$ as the hyperparameter for our main experiments.
Iteration number $N$. This hyper-parameter controls the total number of iterations in the knowledge-guided attack process, akin to training epochs in model optimization. As shown in Figure 5, increasing $N$ leads to improved performance across multiple evaluation metrics (e.g., ASV, RASR, top3_att), indicating that more iterations enable the attack to better manipulate retrieved document stances and influence the final LLM output. For example, ASV improves from 0.56 at $N=1$ to 0.67 at $N=9$, while RASR follows a similar upward trend.
However, higher $N$ also introduces greater computational cost. We observe that performance improves significantly from $N=1$ to $N=5$, but gains plateau between $N=5$ and $N=9$. To balance attack effectiveness and efficiency, we set $N=5$ as the default value in Topic-FlipRAG.
7 Mitigation Analysis and Discussion
To assess the robustness of existing defenses against topic-level opinion manipulation, we evaluate several strategies, including perplexity filtering, random masking, paraphrasing, and reranking. However, our empirical analysis shows that these methods are largely ineffective against Topic-FlipRAG, which leverages subtle, semantics-preserving perturbations that evade conventional detection and intervention. In light of these limitations, we further discuss potential defense directions, including utility-based filtering, TF-IDF anomaly detection, intra-top-$k$ similarity analysis, and certified provable mechanisms grounded in topic and fact consistency.
7.1 Mitigation by Perplexity.
Perplexity (PPL), a standard metric for text quality, is also used to detect adversarial attacks on LLMs and IR models[jain2023baseline, gonen2022demystifying, liu2022order]. High perplexity indicates low-quality or suspicious text. We adopt PPL to identify malicious content.
Figure 6 presents the GPT-2[radford2019language]-evaluated log perplexity distributions of PROCON documents manipulated by various attack methods, highlighting the limitations of perplexity-based defenses. Due to substantial distributional overlap between clean and poisoned documents, setting effective thresholds proves challenging: strict thresholds yield high false positives, while loose ones miss attacks. Among the methods, Collision shows the most noticeable PPL shift, making it relatively detectable. In contrast, Topic-FlipRAG and PAT generate poisoned documents with distributions closely aligned to the original data, posing significant detection challenges. PoisonedRAG, lacking gradient-based optimization, induces negligible PPL changes, further underscoring the insufficiency of perplexity-based filtering for subtle manipulations.
Figure 6: Log perplexity (PPL) distributions on the PROCON dataset.
Figure 7: Attack performance of different random mask rate.
7.2 Mitigation by Random Mask.
@@ 7 Mitigation Analysis and Discussion
Random Mask is a robustness enhancement method that randomly masks a proportion of tokens with a placeholder token $m$ during embedding [zeng2023certified]. By averaging predictions over multiple masked input versions, it reduces sensitivity to small perturbations and enhances model stability. In our implementation, we averaged outputs over three masked copies of each input to assess performance under varying mask rates.
@@
Figure 7 presents experimental results on the PROCON-Dataset targeting the CON stance. When the mask rate is below 0.3, Topic-FlipRAG achieves notably higher manipulation metrics (RASR > 55%, top3-v > 20%) compared to other baselines, underscoring its robust manipulation capability. As the mask rate increases, the effectiveness of all methods declines; however, Topic-FlipRAG still maintains the highest attack success among the baselines. Nonetheless, excessively large mask rates also degrade normal RAG performance, revealing the trade-off between heightened defense and overall system functionality. Consequently, Random Mask alone is insufficient against sophisticated attacks like Topic-FlipRAG, indicating that additional measures are required to strengthen system robustness without undermining functionality.
7.3 Mitigation by Paraphrasing
Cheng et al.[cheng2024trojanrag] and Zou et al. [zou2024poisonedrag] investigate the application of paraphrasing defenses to evaluate the effectiveness of the RAG attacks they proposed. The underlying concept is to modify the query through rewriting, thereby increasing the difficulty for the attacker to successfully execute an attack on a specific query. The prompt used in our approach is provided in Appendix A.4.
The attack effectiveness of Topic-FlipRAG and other baselines against paraphrasing is presented in Table 8. Both the baselines and Topic-FlipRAG exhibit a decline in attack effectiveness when confronted with paraphrasing, yet Topic-FlipRAG remains largely unaffected: its RASR drops marginally from 51.47% to 50.96%, while top3-v decreases from 27.40% to 22.46%. These results suggest that paraphrasing is insufficient to defend against manipulations targeting the topical level. Topic-FlipRAG remains effective because it operates over entire topic-query sets rather than relying on specific query phrasings. Although paraphrasing alters surface wording, it typically preserves the original semantic focus, keeping the rewritten query within the same topic scope. This resilience underscores the difficulty of defending against topic-level RAG manipulations and highlights the need for more advanced, semantics-aware countermeasures.
Figure 8: Comparative analysis of the performance metrics(RASR and top3-v) across different baselines before and after paraphrasing mitigation.
7.4 Mitigation by Reranking.
@@ 7.3 Mitigation by Paraphrasing
We investigate a reranking mechanism within the RAG framework to mitigate adversarial attacks. This approach employs a different IR model to reorder initial retrieval results, thereby reducing attack effectiveness. The defense rationale is twofold: (1) most attacks rely on approximating the target IR model [cho2024typos, liu2022order, liu2024multi], and introducing a distinct reranker disrupts this dependency; (2) different IR models exhibit varying robustness, and reranking enhances overall resilience.
@@
7.5 Future Potential Defense Mechanisms
We have integrated several widely adopted defense strategies from prior work on RAG-based and IR-based attacks, including filtering by perplexity, random masking, paraphrasing, and reranking-based optimization. While these techniques are commonly used, our empirical results demonstrate that they provide only limited robustness against the proposed attack method. To guide the development of more effective future defenses, we outline three promising directions informed by recent insights and attack characteristics:
Filtering via Usefulness Features. Recent RAG-related attack studies reveal that adversaries often enhance the relevance of poisoned documents without improving their actual usefulness in answering user queries. This decoupling creates opportunities for detection. For instance, RbFT [tu2025rbft] highlights the potential of usefulness-based filtering, such as leveraging LLM-based judgments or dedicated retrieval-utility models, to complement relevance-centric ranking systems. Incorporating usefulness estimation may significantly enhance filtering granularity in adversarial settings.
TF-IDF Based Detection. In topic-specific attack scenarios, adversaries frequently inject or substitute high-impact keywords to manipulate semantic relevance, leading to elevated TF-IDF scores. While such high TF-IDF values may also occur in naturally high-ranking documents, the distributional shift introduced by poisoned inputs, especially when paired with usefulness signals, could provide an effective basis for anomaly detection. This hybrid detection approach remains an open but promising research avenue.
Figure 9: Comparative analysis of the performance metrics (RASR and top3-v) across different baselines before and after reranking mitigation.
Intra-Top-k Document Similarity. Under benign retrieval conditions, the top-k documents typically exhibit high semantic coherence, often reflected in strong pairwise similarity. In contrast, targeted poisoning tends to inject documents that diverge semantically from clean results, despite being top-ranked. This reduction in intra-group similarity offers a lightweight and model-agnostic signal that could flag manipulated retrieval outputs without additional supervision. Future work may explore leveraging this pattern for online detection in black-box RAG deployments.
8 Conclusion